A shipment of parts contains 140 items of which 20 are defective. Two of these items
are randomly chosen and inspected. Let X denotes the defective numbers. Find the
probability distribution of X.
1
Expert's answer
2020-12-03T19:56:02-0500

P(x=0)=((140-20)/140)*((139-20)/139)=0.7338

P(x=2)=(20/140)*(19/139)=0.0195

P(x=1)=1-0.0195-0.7338=0.2467